# The CMS Phase-2 Fast Beam Condition Monitor prototype test with beam

### v2

@@ 1.1 Motivation and requirements for the dedicated CMS luminosity detector
During HL-LHC wide range of an average rate of interactions per bunch crossing (also referred to as pileup) must be covered: from up to . Linearity of the luminometer response for the full range and its stability are two key aspects which define the design of a luminometer. Linearity across a wide pileup range is essential as the specialized calibration runs are conducted at a low pileup of a few times per year, while standard physics data taking in proton-proton collisions during HL-LHC operations will occur at . The linearity of FBCM was studied and optimized in simulations [1]. Detector is shown to be linear up to withing uncertainty band. Stability of the detector is defined as ability to maintain its metrological properties, such as accuracy and performance, over extended period of time (on the scale of several years for luminometer). The stability of the luminometer is influenced by operational aspects, as well as the aging of the sensors and electronics under radiation exposure.
@@ 1.2 FBCM detector overview
The FBCM is designed to operate independently of the CMS central trigger and data acquisition systems, functioning continuously with triggerless readout. It employs a dedicated front-end application-specific integrated circuit (ASIC) to amplify signals from CO2-cooled silicon-pad sensors, offering nanosecond-level timing resolution to measure luminosity and beam-induced background. The modular system consists of two half-discs, each containing twelve modules, installed at both ends of the CMS detector, approximately 280 cm from the interaction point. Service electronics are positioned near the outer edge of the half-discs to mitigate radiation-related aging. The electronics design reuses components from the CMS Tracker for power, control, and readout. The FBCM23 ASIC features six channels with adjustable shaping times, enabling fine tuning to match the sensor leakage current. The FBCM mechanics are described in more detail in Subsec. 1.2.1, electronics – in Subsec. 1.2.2, and sensors – in Subsec. 1.2.3.
@@ 1.2.3 FBCM silicon sensors
| S/N after irradiation | 11 | 8 |
| CCE after irradiation at 600 V | 50% | 90% |
@@
Two-pad sensors were previously used in CMS for the BCM1F detector [13] during Run 3 of the LHC and initially operated at bias voltage of 400 V. As they exhibited current instability after exposure to a fluence of about cm-2 1 MeV neutron equivalent, operational bias voltage was lowered to 350 V. This instability [14] manifested itself in the unpredicted rise of the leakage current and eventual trip due to reached set power supply current limit (50 A); it was attributed to the build up of surface current and the formation of a conductive path, which allowed current to flow from the edges of the sensor to the guard ring. Reduction of the bias voltage reduces the frequency of the trips, but did not mitigate it. Further reduction was not possible due to loose off sensors efficiency. To reproduce this behavior under controlled laboratory conditions, the two-pad sensors were tested under an X-ray lamp [9], which generates only surface charge. Over ten hours of X-ray exposure, a slow increase of the current was observed. Intensity of X-ray was selected to match the leakage current close to that observed in operations at the beginning of collisions (10 A). When the X-ray lamp was switched off, a gradual current decrease was observed, lasting about one hour. To verify that the double guard ring structure of the six-pad sensor resolves this issue, the same test was repeated. The results showed that the current through the guard ring and DC-pads remained stable after hours of the surface irradiation, making the six-pad sensor the preferred choice for FBCM construction. However, both types of sensors were thoroughly tested in the proton beam test in July 2024.
@@ 3.1 Time over threshold
Fig. 17 (right) shows the effect of RC3 versus RC6 ASIC settings on the ToT histogram for the same input channel and at a fixed 0.9 fC threshold value. As the shaping time of the signal is different for these two settings, the expected shift of the ToT to higher value for RC6 is observed. This is beneficial for irradiated sensors, as it compensates for the lower charge collection efficiency, but has to be operated in combination with higher thresholds to cut the noise. In this particular example utilizing not irradiated sensor, RC6 integral is about 15% lower than that of RC3, indicating efficiency drop. While it was checked that for the irradiated sensors the integrals of the histograms aggregated for the same number of spills with RC3 and RC6 settings agree within 5%.
@@ 3.2 Time walk
The main deliverable of FBCM for HL-LHC will be ToA, which will be used for bunch-by-bunch luminosity and beam-induced background measurements. The ToA with respect to the bunch crossing will be aggregated in sub-bunch crossing (foreseen to be 2.5 ns) granularity histograms during 1 second time periods. During the beam test ToA is not meaningful in the absence of a time reference. During LHC operation, the beginning of the LHC orbit and the LHC bunch clock will serve as such references.

### v3

@@ 1.1 Motivation and requirements for the dedicated CMS luminosity detector
During the HL-LHC phase, a large range of average rate of interactions per bunch crossing (also referred to as pileup, ) must be covered: from up to . The linearity and the stability of the luminometer response to the full range are two key aspects which define its design. The linearity across a wide pileup range is essential as the specialized calibration runs are conducted at a low pileup of a few times per year, while standard physics data taking in proton-proton collisions during HL-LHC operations will occur at . The linearity of FBCM was studied and optimized in simulations [1]. The detector is shown to be linear up to within the uncertainty band. The stability of the detector is defined as the ability to maintain its metrological properties, such as accuracy and performance, over an extended period of time, on the scale of several years for luminometer. The stability of the luminometer is influenced by operational aspects, as well as the aging of the sensors and electronics under radiation exposure.
@@ 1.2 FBCM detector overview
The FBCM is designed to operate independently of the CMS central trigger and data acquisition systems, functioning continuously with triggerless readout. It employs a dedicated front-end application-specific integrated circuit (ASIC) to amplify signals from CO2-cooled silicon-pad sensors, offering nanosecond-level timing resolution to measure luminosity and beam-induced background. The modular system consists of two half-discs, each containing twelve modules, installed at both ends of the CMS detector, approximately 280 cm from the interaction point. Service electronics are positioned near the outer edge of the half-discs to mitigate radiation-related aging. The electronics design reuses components from the CMS Tracker for power, control, and readout. The FBCM23 ASIC features six channels with adjustable shaping times, enabling fine tuning to match the sensor leakage current. The FBCM mechanics are described in more detail in Subsec. 1.2.1, electronics in Subsec. 1.2.2, and sensors in Subsec. 1.2.3.
@@ 1.2.3 FBCM silicon sensors
| Estimated S/N after irradiation | 11 | 8 |
| Estimated CCE after irradiation at 600 V | 50% | 90% |
@@
Two-pad sensors were previously used in CMS for the BCM1F detector [13] during Run 3 of the LHC and initially operated at bias voltage of 400 V. As they exhibited current instability after exposure to a fluence of about cm-2 1 MeV neutron equivalent, operational bias voltage was lowered to 350 V. This instability [14] manifested itself in the unpredicted rise of the leakage current and eventual trip due to having reached the set power supply current limit (50 A). This was attributed to the build up of surface current and the formation of a conductive path, which allowed current to flow from the edges of the sensor to the guard ring. Reduction of the bias voltage minimizes the frequency of the trips, but does not mitigate them. Further bias voltage reduction was not possible due decline in sensors efficiency. To reproduce this behavior under controlled laboratory conditions, the two-pad sensors were tested under an X-ray lamp [9], which generates only surface charge. Over ten hours of X-ray exposure, a slow increase of the current was observed. Intensity of X-ray was selected to match the leakage current close to that observed in operations at the beginning of collisions (10 A). When the X-ray lamp was switched off, a gradual current decrease was observed, lasting about one hour. To verify that the double guard ring structure of the six-pad sensor resolves this issue, the same test was repeated. The results showed that the current through the guard ring and DC-pads remained stable after hours of the surface irradiation, making the six-pad sensor the preferred choice for FBCM construction. However, both types of sensors were thoroughly tested in the proton beam test in July 2024.
@@ 3.1 Time over threshold
Fig. 17 (right) shows the effect of RC3 versus RC6 ASIC settings on the ToT histogram for the same input channel and at a fixed 0.9 fC threshold value. As the shaping time of the signal is different for these two settings, the expected shift of the ToT to higher value for RC6 is observed. This is beneficial for irradiated sensors, as it compensates for the lower charge collection efficiency, but has to be operated in combination with higher thresholds to cut the noise. In this particular example utilizing a not irradiated sensor, the RC6 integral is about 15% lower than that of RC3, indicating a drop in efficiency. It was checked that for the irradiated sensors the integrals of the histograms aggregated for the same number of spills with RC3 and RC6 settings agree within 5%.
@@ 3.2 Time walk
The main deliverable of the FBCM for HL-LHC will be ToA information, which will be used for bunch-by-bunch luminosity and beam-induced background measurements. The ToA with respect to the bunch crossing will be aggregated in sub-bunch crossing (foreseen to be 2.5 ns) granularity histograms during 1 second time periods. During the beam test ToA is not meaningful in the absence of a time reference. During LHC operation, the beginning of the LHC orbit and the LHC bunch clock will serve as such references.